Ch6 Sec2: Problem 3 PreviousProblem List Next (1 point) A random sample of 100 observations from a population with standard deviation 9.13 yielded a sample mean of 91.6 1. Given that the null hypothesis is--90 and the alternative hypothesis is μ > 90 using a) Test statistic- (b) P- value: (c) The conclusion for this test is: 05, find the following: A. Reject the null hypothesis B. There is insufficient evidence to reject the null hypothesis C. None of the above 2. Given that the null hypothesis is μ-90 and the alternative hypothesis is μ (a) Test statistic (b) P -value: (c) The conclusion for this test is: 90 using a-.05, find the following: A. There is insufficient evidence to reject the null hypothesis B. Reject the null hypothesis C. None of the above
Answer #1

Here

sample mean ar{x}= 91.6

population standard deviation σ 9.13

and sample size n 100

The test statistic can be written as

/12 (Z-90 which under H0 follows a standard normal distribution.

We reject H0 at 5% level of significance if P-value < 0.05

Now,

The value of the test statistic = ~obs 1.75246

and P-value = P(Z > zobs) = P(Z > 1.75246) 1-0.960153 = 0.0398470

Since P-value < 0.05, so we reject the null hypothesis.

2. The value of the test statistic = ~obs 1.75246

and P-value = P(Zobs2 P(Z1.75246) 2(1-0.960153) 20.0398470 0.079694

Since P-value > 0.05, so

ans-> A) There is insufficient evidence to reject the null hypothesis.